OpenStudy (anonymous):

Solve for x. (x-10) + (2x+1)=90

OpenStudy (anonymous):

x=13.268 x=3.768

OpenStudy (anonymous):

\[2x+1+x-10=90\] 3x-9=90 3x=90+9 3x=99 x=33